OpenStudy (anonymous):

hey @Phebe Solve. n + (–14) = 9 A. –23 B. –5 C. 5 D. 23

Nnesha (nnesha):

try to solve this one

OpenStudy (phebe):

hey lol

OpenStudy (phebe):

need anything?????? :D

OpenStudy (phebe):

ohh nvm i see ur Q

Nnesha (nnesha):

okay try this one first and let us know what you get

Nnesha (nnesha):

same like last two

OpenStudy (phebe):

n+(14)-(9)=0 Step by step solution : Step 1 : Equation at the end of step 1 : Step 2 : Solve n+5 = 0 Solving a Single Variable Equation : 2.1 Solve : n+5 = 0 Subtract 5 from both sides of the equation : n = -5 ANSWER IS: n = -5

OpenStudy (anonymous):

5

Nnesha (nnesha):

at least let him/her solve this first

OpenStudy (phebe):

Answer is B-5

OpenStudy (anonymous):

That's not right Phoebe plugging in -5 for n

OpenStudy (phebe):

how its not right?

Nnesha (nnesha):

 first distribute